maria's father is now twice as old as she is, 16 years ago he was four times as old. how old is each now
Answer by Cromlix(4381)

Hi there,
maria's father is now twice as old as she is.
F = 2M......(1)
16 years ago he was four times as old.
F - 16 = 4(M - 16)......(2)
F - 16 = 4M - 64........(2)
Replace F in Eq(2) with 2M
from Eq(1)
2M - 16 = 4M - 64
Collect like terms
2M - 4M = - 64 + 16
-2M = -48
M = -48/-2
M = 24.
Maria is 24 years old
Father is 48 years old.
Hope this helps :-)